You are conducting a study to see if the accuracy rate for fingerprint identification is significantly more than 79%. With Ha : p >79% you obtain a test statistic of z=2.78
. Find the p-value accurate to 3 decimal places.

Answer by ewatrrr(24785)


Hi
Ho:  p = .79
Ha:  p > .79
 test statistic: z = 2.78
p-value = P(z > 2.78)= 1 - P(z ≤  2.78) = 1 - .9973 = .0027 0r .003
Wish You the Best in your Studies.